OpenStudy (anonymous):

A car starts from rest and accelerates at a rate of 3.00 m/s^2 for 10 seconds and then maintains a constant speed for 25 seconds. The car then decelerates at -2.50 m/s^2 until it stops. Find the total distance covered.

OpenStudy (anonymous):

its too hard :( im looking for examples

OpenStudy (anonymous):

PHYSICS!!!!!

OpenStudy (anonymous):

WHY DOES NOBODY HELP ME

OpenStudy (anonymous):

Lol warood try giving dominique some help if mines to hard

OpenStudy (anonymous):

THANK YOU SO MUCH DWADE

OpenStudy (anonymous):

BTW DWAYNE WADE IS A FRKIN BEAST

OpenStudy (anonymous):

HELL YEA! i hate the heat though just wade is cool

OpenStudy (anonymous):

wade has always been cool

OpenStudy (anonymous):

It's three parts, right? 

OpenStudy (anonymous):

D = d1+d2+d3

OpenStudy (anonymous):

Total Distance = DT DT = D1 + D2 + D3 (D1/2/3 are different parts of acceleration) D1 = vi*t + 1/2at^2 D1 = 0 + .5*(3m/s^2)*10^2 D1 = figure it out ;-) D2 = vi*t + 1/2at^2 vi of D2 = 3m/s^2 * 10s = 30m/s D2 = 30m/s*25 + 0 (acceleration = 0) D3 = vi*t +1/2at^2 D3 = 30*t + 1/2*(-2.5m/s^2)*t^2 t = 30/2.5 = 12 D3 = 30*12 + 1/2 (-2.5m/s^2)*30^2 Plug 'n chug

OpenStudy (anonymous):

so Vi is the initial velocity?

OpenStudy (anonymous):

Vi is the initial velocity *in that section* So Vi in the first part is different from Vi in the 2nd and 3rd part.

OpenStudy (anonymous):

yah... sorry for not delineating that. I have never typed physics before :D

OpenStudy (anonymous):

yea its weird its not your fault :P

OpenStudy (anonymous):

:) no worries

OpenStudy (anonymous):

so let me get this straight cuz im an idiot when it comes to physics, for the d1 i got 59 is that remotely correct?

OpenStudy (anonymous):

I don't think so. What did you use for vi = ? t = ? a = ?

OpenStudy (anonymous):

good luck guys... gotta go write a paper -.-

OpenStudy (anonymous):

for vi i used 3.00 for t i used 10 and for a used .5?

OpenStudy (anonymous):

alright @micah.daniel.anderson thanks for the help

OpenStudy (anonymous):

"A car starts from rest and accelerates at a rate of 3.00 m/s^2 for 10 seconds " \[d_1 = v_{i}^2 + \frac{1}{2}at^2\] So how fast is the car going in the beginning?

OpenStudy (anonymous):

does that say v^2?

OpenStudy (anonymous):

would it be 0+1/5*3.00*10^2?

OpenStudy (anonymous):

i mean 0+1/2*3.00*10^2

OpenStudy (anonymous):

is the car moving at 45 m/s in the beginning?

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

oh , nononon

OpenStudy (anonymous):

you were right before, with Vi = 0

OpenStudy (anonymous):

and that's (Vi)^2, yeah

OpenStudy (anonymous):

oh alright

OpenStudy (anonymous):

and so a = 3.00 and t = 10 correct?

OpenStudy (anonymous):

yes yes :)

OpenStudy (anonymous):

so then the equation would read 0^2+0.5*3.00*10^2?

OpenStudy (anonymous):

yup yup - just remember all your units if you're writing it on paper :) (they get messy when you type em, I know :P)

OpenStudy (anonymous):

ill keep that in mind lol. so the outcome gives me 150 thats meters per second?

OpenStudy (anonymous):

or is that just the distance covered initially?

OpenStudy (anonymous):

You used this equation, right? \[d = v_i t + \frac{1}{2}at^2\] *distance* is equal to velocity times time plus.... so units of distance in this case are meters, m It's easy to remember units when you say what the terms mean as opposed to just letters :) And yeah, that's the distance covered in this section: